# Chan-wei Hu1, Chengxuan Qian1, Huaxiu Yao4, Zhengzhong Tu [^1] footnotemark: 1

### v2

@@ 7 Conclusion
Although Re-Align has demonstrated superior performance on both hallucination and general VQA benchmarks, it does not always achieve state-of-the-art results on general tasks; in some cases, its performance is even worse than that of vanilla VLMs. Future research could explore strategies to eliminate this alignment tax or identify an optimal balance for this trade-off. The potential risks of this work align with the general challenges of RLHF alignment. As more powerful alignment techniques are developed, they may inadvertently empower adversarial approaches that exploit these models, potentially leading to unfair or discriminatory outputs. Meanwhile, these adversarial strategies can be used to generate negative samples, which can ultimately contribute to the development of more robust and reliable VLMs over time.
@@ Appendix A Overview of Re-Align
1:$\mathcal{D}\leftarrow\emptyset$ // Init preference dataset
2:$N\leftarrow|\{v_{i}\}|$
@@
19:Update $\pi_{\theta}$ through $\mathcal{L}_{\text{rDPO}}$ (eq. 1)
@@ Appendix D Examples of Preference Pair
Table 4 and 5 provide examples of the constructed preference data for the VQA and image captioning, and each data sample contains textual instruction, input image, retrieved image, chosen response, and rejected response.
@@ Appendix E Response Examples
Appendix F Additional Discussions
Performance Variations on General VQA tasks
Segment-level Preference
Building on the findings of Yu et al. (2024b), we generate preference data by inducing hallucinations at the segment level than at the sentence level (as seen in approaches such as POVID (Zhou et al., 2024a), STIC (Deng et al., 2024), and CSR (Zhou et al., 2024b)), to provide robust supervision signals during the alignment process. This finer-grained preference modeling yields clearer and more precise learning signals, enabling the model to better distinguish between subtle hallucinations and ground truth responses. To further investigate these segment-level preference signals, we expanded the finetuning dataset from $11k$ to $16k$ image samples. As illustrated in Figure 7, when using LLaVA-v1.6-Mistral-7B as the backbone with ScienceQA as the case study, Re-Align achieved a significant performance improvement—from $0.45$ to $1.34$—demonstrating the effectiveness of our approach.
@@ Appendix G Licenses
For the general VQA benchmarks, ScienceQA (Lu et al., 2022), TextVQA (Singh et al., 2019), MM-Vet (Yu et al., 2023b), VisWiz (Gurari et al., 2018), LLaVABench (Liu, 2023), and MMBench (Liu et al., 2024c) are released under MIT, CC BY 4.0, Apache-2.0, CC BY 4.0, Apache-2.0, and Apache-2.0 licenses respectively. While MME (Fu et al., 2023) was released without an accompanying license.
@@ Appendix I Computational Cost
All finetuning and evaluation experiments were executed on four NVIDIA A6000ada GPUs. Table 7 details the time required for Re-Align to fine-tune each model.
@@
LLaVA-v1.6-Vicuna-7B 46 min
LLaVA-v1.6- Vicuna-13B 72 min
Table 7: Time required for finetuning VLMs with Re-Align.
@@
For all the experiments, we finetuning VLMs with Re-Align for 1 epoch. We deploy LoRA finetuning with lora_r=128, lora_alpha=256, target_module=all, and hyperparameters as presented in Table 8.
@@
Table 8: Hypeterparameter setting for finetuning.
@@
  • Democratizing access to robust AI: Our method can democratize access to advanced mutimodal AI models under low-resource or data-scarce settings, which empowers researchers and practitioners with limited computational resources to participate in cutting-edge AI development, ultimately contributing to a more equitable and diverse AI ecosystem.

### v3

@@ 7 Conclusion
Although Re-Align has demonstrated superior performance on both hallucination and general VQA benchmarks, it does not always achieve state-of-the-art results on general tasks; in some cases, its performance is even worse than that of vanilla VLMs. Future research could explore strategies to eliminate this alignment tax or identify an optimal balance for this trade-off.
The potential risks of this work align with the general challenges of RLHF alignment. As more powerful alignment techniques are developed, they may inadvertently empower adversarial approaches that exploit these models, potentially leading to unfair or discriminatory outputs. Meanwhile, these adversarial strategies can be used to generate negative samples, which can ultimately contribute to the development of more robust and reliable VLMs.
@@ Appendix A Overview of Re-Align
[…]
@@
19:Update $\pi_{\theta}$ through $\mathcal{L}_{\text{rDPO}}$ (eq.˜1)
@@ Appendix D Examples of Preference Pair
Table 5 and 6 provide examples of the constructed preference data for the VQA and image captioning, and each data sample contains textual instruction, input image, retrieved image, chosen response, and rejected response.
@@ Appendix E Response Examples
Appendix F Data Curation
Table 8 summarizes the key characteristics of the preference datasets employed by Re-Align and several baseline alignment methods. Importantly, the reported dataset sizes correspond only to the preference pairs used directly for alignment training, and not to the total datasets leveraged in each pipeline. Several baseline methods, such as LLaVA-RLHF and POVID, additionally rely on larger supervised fine-tuning stages with external datasets, whereas Re-Align operates solely on curated preference data.
Unlike baselines that depend on synthetic perturbations or expensive human annotations, Re-Align introduces a semantically-guided image retrieval and masking procedure to construct preference datasets. This strategy offers several critical advantages:
  • Semantic Coherence. Retrieved natural images preserve contextual integrity and semantic relationships, which are often degraded by cropped or artificially edited images.
  • Natural Preference Signals. The curated pairs reflect genuine visual understanding rather than superficial low-level perturbations (e.g., Gaussian noise, color jitter, or downsampling artifacts).
The construction of preference data is a key determinant of downstream alignment performance. By leveraging semantically-guided retrieval, Re-Align produces preference pairs that are both semantically rich and visually natural, contributing to its robustness across both general VQA and hallucination benchmarks.
@@ Appendix G Licenses
For the general VQA benchmarks, ScienceQA (Lu et al., 2022), TextVQA (Singh et al., 2019), MM-Vet (Yu et al., 2023b), VisWiz (Gurari et al., 2018), LLaVABench (Liu, 2023), and MMBench (Liu et al., 2024d) are released under MIT, CC BY 4.0, Apache-2.0, CC BY 4.0, Apache-2.0, and Apache-2.0 licenses respectively. While MME (Fu et al., 2023) was released without an accompanying license.
@@ Appendix I Computational Cost
All fine-tuning and evaluation experiments were executed on four NVIDIA A6000ada GPUs. Table 9 details the time required for Re-Align to fine-tune each model.
@@
LLaVA-v1.6-Vicuna-7B
46 min
LLaVA-v1.6- Vicuna-13B
72 min
Table 9: Time required for fine-tuning VLMs with Re-Align.
@@
For all the experiments, we fine-tuning VLMs with Re-Align for 1 epoch. We deploy LoRA fine-tuning with lora_r=128, lora_alpha=256, target_module=all, and hyperparameters as presented in Table 10.
@@
Table 10: Hypeterparameter setting for fine-tuning.
@@
  • Democratizing access to robust AI: Our method can democratize access to advanced multimodal AI models under low-resource or data-scarce settings, which empowers researchers and practitioners with limited computational resources to participate in cutting-edge AI development, ultimately contributing to a more equitable and diverse AI ecosystem.
Appendix L Broader Impacts
The research presented in this paper, particularly the development of the Re-Align framework, has significant broader impacts that extend beyond the immediate technical contributions. By improving the alignment of Vision Language Models (VLMs), our work contributes to the creation of more reliable, trustworthy, and capable AI systems, which have profound implications for various societal domains.
A primary impact of this research is the enhancement of safety and trustworthiness in AI systems deployed in critical applications. The reduction of hallucinations is paramount for autonomous systems where perception and decision-making must be grounded in reality. For instance, in autonomous driving, reliable visual understanding is non-negotiable. Our work aligns with efforts to build end-to-end autonomous driving models (Xing et al., 2025b; Luo et al., 2025), improve motion prediction through equivariant geometry (Wang and Chen, 2023b, a), and multi-agent communication (Wang et al., 2025c, a). By ensuring that a VLM’s outputs are faithful to its visual inputs, Re-Align contributes to the foundational safety required for deploying these technologies. The principles extend to other domains like robotics and collaborative agent systems, where trustworthy AI is essential for safe and effective operation (Li et al., 2025a; Gao et al., 2025a; Chen et al., 2024a).
Furthermore, our work contributes to the broader unification and advancement of generative and discriminative AI models. The alignment techniques we propose are part of a larger trend towards creating more cohesive and capable foundation models (Liu et al., 2024c). This advancement enables a wide range of new applications. For example, improved visual fidelity is crucial for tasks like novel view synthesis from single RGBD images (Hetang and Wang, 2023) and for understanding complex 3D environments from partial data (Zhang et al., 2021). As these models become more robust, they can be applied to creative industries, virtual reality, and scientific visualization with greater confidence.
Finally, the development of more effective and efficient alignment techniques has implications for the accessibility and democratization of AI. As methods like Direct Preference Optimization (DPO) become more refined, they can potentially lower the barrier to fine-tuning powerful models for specific, beneficial purposes. Techniques that improve the learning process, such as prompt learning using metaheuristics (Pan et al., 2024), can make the customization of large models more efficient. However, it is crucial to acknowledge the dual-use nature of these powerful technologies. The same methods that align models to be helpful and harmless could potentially be used for malicious purposes. Therefore, ongoing research into robust safety protocols, ethical guidelines, and trustworthiness benchmarks (Xing et al., 2024) is essential to mitigate these risks and ensure that the societal benefits of advanced AI systems like those improved by Re-Align are realized responsibly.